# Missing Pieces: How Framing Uncertainty Impacts Longitudinal Trust in AI Decision Aids — A Gig Driver Case Study

### v1

@@ 6.1.3. Participants
F1: A 47-year-old male with a high school degree who drives for Instacart and Lyft in Los Angeles, Condition 2
@@
F2: A 42-year-old female with a high school degree who drives for Uber Eats in Los Angeles, Condition 2
@@
F3: A 29-year-old female with a graduate degree who drives for Lyft in Chicago, Condition 4
@@
F4: A 49-year-old male with an undergraduate degree who drives for Lyft and Uber in Los Angeles, Condition 4
@@
F5: A 46-year-old male with an undergraduate degree who drives for DoorDash, GrubHub, and Uber Eats in Chicago, Condition 3
@@
F6: A 39-year-old male with a graduate degree who drives for Lyft, Uber, and Uber Eats in Los Angeles, Condition 2
@@
F7: A 39-year-old male with a professional degree who delivers for DoorDash in Los Angeles by bike, Condition 4
6.1.4. Analysis
For quantitative analysis, we applied statistical models implemented in statsmodels 0.14.0 (Seabold and Perktold, 2010) to our measurements. The questions comprising the instruments were presented to participants as 5-point Likert-type scales in ascending order from 1–5, following (Chyung et al., 2017, 2018). Ordinary least squares (OLS) models were fitted to pre- and post-interaction (Survey 2 and Survey 4) trust scores. For Survey 2 trust scores, we took the average of the 5 items drawn from (Madsen and Gregor, 2000); for the Survey 4 trust score, we first inverted 2 items measuring distrust from (Jian et al., 2000) and then took their average with 3 other items from (Jian et al., 2000). To model intra-subject dependencies, linear mixed models (LMMs) were fitted to daily data from Survey 3, with participant IDs as random effects. For the trust score, we took the average of the 3 items drawn from (Madsen and Gregor, 2000). We also computed a reliance indicator, which was 0 if a participant (a) indicated their desire to decrease or pause their use of the tool on the following day or (b) failed to complete the following day’s Survey 3, and 1 otherwise. For each model, we started with all relevant variables, and then manually performed variable selection to minimise the Akaike information criterion (AIC) (Cavanaugh and Neath, 2019). The reliance indicator is more properly fit by a logistic generalised linear mixed model (GLMM). However, since the statsmodels implementation of GLMMs did not support either -values or the AIC, model selection was performed using the LMM equivalent. We report results for both.
We used the same methodology as the pilot interviews (Section 5.2) to analyse the interviews for the longitudinal trials.
6.2. Results
Of the 51 participants, 33 went on to complete all 7 days of interaction with the tool, thus receiving $40 of compensation. We show full retention statistics for each of the 4 conditions in Figure 2. Overall, although individual participants differed in their level of reliance on the recommendations (see also Section 6.2.2), participants generally felt that our estimates were reasonably calibrated. Across all daily interactions corresponding to Survey 3 responses, participants reported that their earnings for the day felt roughly equal to the tool’s estimate after 49% of interactions, higher than the tool’s estimate after 23% of interactions, and lower than the tool’s estimate after 28% of interactions — a roughly symmetric distribution. Nevertheless, trust and reliance still significantly differed between conditions, as is also evident from the differences in retention rates in Figure 2. We now explore these differences through our quantitative results.
6.2.1. Quantitative Results
In describing the statistical models that we fitted, we use the following variable notations: for the Survey 2 trust score; for the Survey 3 trust score on day ; for the Survey 4 trust score; for the reliance indicator on day ; for the participant’s schedule page condition; for the monetary difference between a participant’s actual and estimated earnings for the hours that they worked on day ; for the monetary difference between a participant’s actual and estimated earnings for the hours that the tool recommended on day ; for an indicator denoting whether the participant perceived their earnings to be at least as high as the tool’s estimate; for an indicator denoting whether the participant perceived their earnings to be at least as high as their expectations; and as the day for a data point from Survey 3. Statistically significant coefficients are denoted using asterisks for each significance level: * (0.05), ** (0.01), *** (0.001).
Research Question 1
This research question assesses the relationship between users’ perceived accuracy of the tool’s estimates and trust. As applied to our context, Hypothesis 1 hypothesises that trust would be increased by users perceiving their earnings as being at least as good as the tool’s estimates. Based on the trust scores self-reported during and after tool interactions (Surveys 3 and 4), we consider Hypothesis 1 to be partially supported.
For the daily Survey 3 trust score, the best-fitting LMM had as fixed effects, and thus depended only on interactions of the day with (1) the monetary difference in earnings for the participant’s worked hours and (2) the perceived accuracy of the tool’s estimate. When participants perceived their earnings to be at least as high as the tool’s estimates, their trust scores were significantly higher (; vs ). We also observed significant longitudinal effects: the trust score generally increased with each passing day, but the rate of increase was significantly lessened among daily interactions where participants perceived the tool’s estimates as being accurate (; vs ). This result can likely be attributed to the higher baseline level of trust for participants who perceived the tool’s estimates as being accurate. Fitted coefficients are shown in Table 1.
Coefficient | |||
Fixed effects | |||
Intercept | 3.158 | 0.233 | 0.000*** |
0.133 | 0.044 | 0.003** | |
0.004 | 0.002 | 0.004** | |
-0.001 | 0.000 | 0.010** | |
0.543 | 0.233 | 0.020* | |
-0.122 | 0.050 | 0.014* | |
Random effect | 0.527 | 0.258 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 trust. Significant coefficients are the intercept (3.158), (0.133), (0.004), the interaction of and (-0.001), (0.543), and the interaction of and (-0.122).
However, we failed to find significant evidence that the outcomes of individual interactions improved participants’ trust when considering their experiences retrospectively. For the post-interaction Survey 4 trust score, the best-fitting OLS model was simply , and thus only depended on the perceived accuracy of the tool’s estimate for the first day. This variable did not have a significant effect (; vs ). Still, the positive correlation of this variable with post-interaction trust echoes the findings from the longitudinal studies of (Mou and Cohen, 2017; Mou et al., 2017) regarding the importance of initial experiences. Fitted coefficients are shown in Table 2.
Coefficient | |||
Intercept | 3.3600 | 0.292 | 0.000*** |
0.4900 | 0.317 | 0.132 |
Table of ordinary least squares model coefficients for post-interaction Survey 4 trust. The intercept (3.36) is significant, but the coefficient for (0.49) is not.
Research Question 2
This research question assesses the relationship between users’ perceived accuracy of the tool’s estimates, trust, and reliance. As applied to our context, Hypothesis 2 hypothesises that, holding trust constant, reliance would be increased by users perceiving their earnings as being at least as good as the tool’s estimate. Based on the daily Survey 3, we consider Hypothesis 2 to be supported. For the Survey 3 reliance indicator (whether participants increased or maintained their intended level of reliance), the best-fitting LMM that included the trust score had as fixed effects, and thus only depended on (1) the perceived accuracy of the tool’s estimate, (2) the perceived accuracy of the participant’s expectations, and (3) the trust score. Removing the trust score improved the AIC; however, we retained it to isolate the effect of perceived accuracy on reliance independent of trust. Holding the trust score constant, we found significant evidence that reliance was higher when participants perceived the tool’s estimates as being accurate (; LMM means vs ; GLMM log-odds vs ). Fitted coefficients for both the LMM and GLMM are shown in Table 3. However, -values for the GLMM could not be computed.
LMM | GLMM | ||||
Coefficient | |||||
Fixed effects | |||||
Intercept | 0.411 | 0.140 | 0.003** | -1.140 | 0.6619 |
0.214 | 0.073 | 0.003** | 1.530 | 0.457 | |
0.069 | 0.066 | 0.299 | -0.066 | 0.465 | |
0.029 | 0.033 | 0.389 | 0.412 | 0.178 | |
Random effect | 0.074 | 0.088 | -43.000 | 0.181 |
Table of linear mixed model and generalised linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 reliance. Significant coefficients are the intercept (0.411 for the LMM, -1.14 for the GLMM) and (0.214 for the LMM, 1.53 for the GLMM), where significance is based on the LMM -values only.
Research Question 3
This research question assesses the effect of framings of uncertainty on trust and reliance by comparing them across the different conditions outlined in Section 4.2. Relative to the baseline of Condition 1, Hypothesis 3 hypothesises that designs which present estimates with higher granularity but not error visibility (Condition 2) increase trust and reliance, while Hypothesis 4 hypothesises that designs which present ranges for estimates (Conditions 3 and 4) decrease trust and reliance. Based on the trust scores and reliance indicators from all surveys, we consider Hypothesis 3 and Hypothesis 4 to be partially supported.
For the pre-interaction Survey 2 trust score, the best-fitting OLS model was , and thus depended only on the condition. Group means and pairwise comparisons are shown in Table 4. Pairwise comparison means are expressed as the mean of the second group minus the mean of the first.
Coefficient | |||
Group means | |||
Condition 1 | 3.339 | 0.266 | 0.000*** |
Condition 2 | 3.629 | 0.369 | 0.000*** |
Condition 3 | 4.183 | 0.384 | 0.000*** |
Condition 4 | 3.367 | 0.384 | 0.000*** |
Pairwise comparisons | |||
1 - 2 | 0.290 | 0.369 | 0.432 |
1 - 3 | 0.845 | 0.384 | 0.028* |
1 - 4 | 0.028 | 0.384 | 0.941 |
2 - 3 | 0.555 | 0.377 | 0.141 |
2 - 4 | -0.262 | 0.377 | 0.487 |
3 - 4 | -0.817 | 0.391 | 0.037* |
Table of pairwise comparisons for the ordinary least squares model for pre-interaction Survey 2 trust. The group means for the four conditions are 3.339, 3.629, 4.183, and 3.367. Significant differences exist in Conditions 1 and 3 (difference 0.845) and Conditions 3 and 4 (difference -0.817).
For the daily Survey 3 trust score, the best-fitting LMM including the condition had as fixed effects, and thus depended only on (1) the condition and interactions of the day with (2) the monetary difference in earnings for the participant’s worked hours and (3) the perceived accuracy of the tool’s estimate. Note the lack of interactions for the condition. Group means, pairwise comparisons, and other coefficients are shown in Table 5.
Coefficient | |||
Group means | |||
Condition 1 | 3.160 | 0.316 | 0.000*** |
Condition 2 | 3.083 | 0.378 | 0.000*** |
Condition 3 | 3.520 | 0.385 | 0.000*** |
Condition 4 | 2.770 | 0.381 | 0.000*** |
Pairwise comparisons | |||
1 - 2 | -0.077 | 0.315 | 0.807 |
1 - 3 | 0.361 | 0.332 | 0.278 |
1 - 4 | -0.389 | 0.319 | 0.223 |
2 - 3 | 0.437 | 0.314 | 0.163 |
2 - 4 | -0.313 | 0.300 | 0.298 |
3 - 4 | -0.750 | 0.316 | 0.018* |
Other fixed effects | |||
0.141 | 0.044 | 0.001*** | |
0.004 | 0.002 | 0.004** | |
-0.001 | 0.000 | 0.011* | |
0.585 | 0.233 | 0.012* | |
-0.130 | 0.050 | 0.009** | |
Random effect | 0.458 | 0.228 |
Table of pairwise comparisons for the linear mixed model for daily Survey 3 trust. The group means for the four conditions are 3.16, 3.083, 3.52, and 2.77. Significant differences exist in Conditions 3 and 4 (difference -0.75). Other significant fixed effects are (0.141), (0.004), the interaction of and (-0.001), (0.585), and the interaction of and (-0.13).
For the daily Survey 3 reliance indicator, the best-fitting LMM including the condition had as fixed effects, and thus only depended on (1) the condition, (2) the perceived accuracy of the tool’s estimate, and (3) the perceived accuracy of the participant’s expectations. We did not include the trust score in this model, as it harmed the AIC and we found previously that it had no significant effect. Note the lack of interactions involving the condition. Group means, pairwise comparisons, and other coefficients for both the LMM and GLMM are shown in Table 6.
LMM | GLMM | ||||
Coefficient | |||||
Group means | |||||
Condition 1 | 0.324 | 0.111 | 0.003** | -0.476 | 0.421 |
Condition 2 | 0.595 | 0.140 | 0.000*** | 0.691 | 0.557 |
Condition 3 | 0.503 | 0.147 | 0.001*** | 0.333 | 0.574 |
Condition 4 | 0.594 | 0.141 | 0.000*** | 0.442 | 0.533 |
Pairwise comparisons | |||||
1 - 2 | 0.272 | 0.130 | 0.037* | 1.167 | 0.484 |
1 - 3 | 0.179 | 0.137 | 0.191 | 0.809 | 0.497 |
1 - 4 | 0.271 | 0.132 | 0.041* | 0.917 | 0.468 |
1 - 3 | -0.092 | 0.129 | 0.473 | -0.358 | 0.529 |
2 - 4 | -0.001 | 0.123 | 0.992 | -0.249 | 0.507 |
3 - 4 | 0.091 | 0.131 | 0.486 | 0.108 | 0.521 |
Other fixed effects | |||||
0.217 | 0.073 | 0.003** | |||
0.070 | 0.066 | 0.293 | |||
Random effect | 0.068 | 0.079 | -43.000 | 0.013 |
Table of pairwise comparisons for the linear mixed model and generalised linear mixed model for daily Survey 3 reliance. The group means for the four conditions are 0.324, 0.595, 0.503, and 0.594 for the LMM, and -0.476, 0.691, 0.333, and 0.442 for the GLMM. For the LMM, significant differences exist in Conditions 1 and 2 (difference 0.272) and Conditions 1 and 4 (difference 0.271). The fixed effect for (0.217) is also significant.
For the post-interaction Survey 4 trust score, the best-fitting OLS model including the condition was , and thus only depended on (1) the condition and (2) the perceived accuracy of the tool’s estimate on the first day. While we found no significant differences in trust scores between conditions, Condition 1 had higher trust scores than Condition 2, 3, and 4. Group means, pairwise comparisons, and other coefficients are shown in Table 7.
Coefficient | |||
Group means | |||
Condition 1 | 3.586 | 0.384 | 0.000*** |
Condition 2 | 3.273 | 0.445 | 0.000*** |
Condition 3 | 3.519 | 0.470 | 0.000*** |
Condition 4 | 3.211 | 0.432 | 0.000*** |
Pairwise comparisons | |||
1 - 2 | -0.313 | 0.345 | 0.364 |
1 - 3 | -0.067 | 0.385 | 0.862 |
1 - 4 | -0.374 | 0.338 | 0.268 |
2 - 3 | 0.246 | 0.345 | 0.475 |
2 - 4 | -0.061 | 0.292 | 0.834 |
3 - 4 | -0.308 | 0.338 | 0.363 |
Other coefficients | |||
0.4972 | 0.325 | 0.137 |
Table of pairwise comparisons for the ordinary least squares model for post-interaction Survey 4 trust. The group means for the four conditions are 3.586, 3.273, 3.519, and 3.211. Significant differences do not exist.
Our hypotheses did not capture the reality that different framings of uncertainty do not uniformly impact trust and reliance. Based on the daily Survey 3 measurements, Conditions 2 and 4 had the lowest reported trust ( for each condition: ) yet significantly higher reliance than Condition 1 ( for pairwise comparisons; for each condition: ). This observation can likely be attributed to the fact that we observed much less attrition in these conditions. It is possible that this was the result of survivorship bias, where participants in Conditions 1 and 3 chose to drop out instead of reporting lower trust scores. Yet, based on responses for Day 1 (when dropout rates were still relatively uniform), Condition 1 and 3 began with higher trust scores ( for each condition: ). Thus, adding either granularity or hedging to estimates may have helped participants to maintain interest in relying on the tool for their decision-making, even if their baseline trust was lower. Such designs may have augmented reliance by calibrating perceptions of its trustworthiness to be realistic but not necessarily optimal (Ferrario and Loi, 2022; Schemmer et al., 2023).
With respect to the effect of ranged estimates, Condition 3 resulted in a significantly higher pre-interaction Survey 2 trust score than Condition 1 (, for Conditions 1/3). This improvement is contrary to the prediction of Hypothesis 4, and was likely due to participants perceiving the ranges as adding useful information at least initially (Section 6.2.2). Compared to the findings of (Prabhudesai et al., 2023), who suggested that presenting uncertainty mitigates undue reliance, we hypothesise that the more task-aligned presentation of uncertainty in our context (direct estimates that abstract away probabilities) led to participants trusting the tool more. While Condition 3’s daily Survey 3 trust score was also higher than Condition 1, the difference was not significant (; for Conditions 1/3), and Condition 3’s post-interaction Survey 4 trust score was lower (but not significantly) than Condition 1 (; for Conditions 1/3). We also note that Condition 3’s trust scores were significantly higher than Condition 4 for both Survey 2 (; for Conditions 3/4) and Survey 3 (; for Conditions 3/4). Hedging may have increased error visibility to the extent that it offset the utility of the ranges, leading participants to think more critically about the estimated earnings and recommendations (as found by (Prabhudesai et al., 2023)) and thus decreasing trust according to Hypothesis 4.
Confounding Variables
Beyond the analyses described above, we performed additional analyses to assess the effect of possible confounding variables on our measurements. To do so, we added variables to the best-fitting models for these measurements that represented the constraint variant (maximum hours or minimum earnings/daily or weekly), the participant’s city (Los Angeles, New York, Chicago, or Houston), and the participant’s level of motivation (as measured by Question 5 in the formative Survey 1, “When I drive, it’s important to me that I make some minimum amount of money”; see Section A.1). We mostly did not find significant differences in trust and reliance measures between the different levels of these variables. There were two exceptions. First, participants who set maximum hours per day had significantly higher reliance for the daily Survey 3 (). We believe that this result can be partially attributed to the smaller number of participants for this variant, as well as one participant who consistently indicated their desire to increase their reliance on the tool every day. Second, participants who were more motivated to make a minimum amount of money had significantly higher trust for the pre-interaction Survey 2 (), but we did not observe any longitudinal effects for this variable. Our full quantitative results for these analyses are detailed in Appendix C.
6.2.2. Qualitative Results
In the following, we denote our interview participants as F1–F7 following the ordering in Section 6.1.3, where we also list full demographic details. Again, F1, F2, and F6 were placed in Condition 2; F5 was placed in Condition 3; and F3, F4, and F7 were placed in Condition 4.
Motivations and Routines
Participants reported a diversity of motivations and routines. F1 and F2 viewed gig driving as a primary source of income, while F3–F6 used it to supplement other sources of income. F1, F4, and F6 maintained consistent driving schedules; F3 and F5 scheduled their driving around other commitments; and F2 and F6 adjusted their schedules based on the level of demand and their own motivations. Meanwhile, F7 differed from other participants in that he only performed deliveries only during his free time and using a bicycle instead of driving. He felt that the tool did not capture the effects of these habits on his potential earnings.
Tool Benefits and Concerns
Except for F4 and F7, all participants felt that the time they spent interacting with the tool was worthwhile. F1–F3, F5, and F6 felt that the tool helped to visualise information about potential earnings in a helpful manner. This addresses a key pain point of gig drivers: F1–F6 had all reported challenges in estimating their potential earnings as a consequence of unpredictability in gig demand, pay, or location, or of gig platforms providing insufficient information. F2, F4, and F6 had also indicated that their earnings were impacted by the variability of bonuses and tips. In terms of other benefits, F1, F2, F5, and F6 felt that the tool provided more or different information, either relative to gig platforms or to Gridwise, and F5 and F6 found that the tool helped them to decide when to work. F4 was unconvinced of the utility of the tool due to the many factors that could influence earnings.
Constraint Design
In terms of alignment, F1, F3, F5, and F6 all felt that the constraints the tool allowed them to input mostly aligned with their goals and routines. F5 and F6 particularly appreciated the hourly granularity of the availability constraints. However, F1 and F3 surfaced a previously unknown design issue: it was not intuitive to them that they could select their availability individually for each time slot, and thus the constraints they had entered were suboptimal. Participants also exhibited perceptions of control over the tool; F1–F3, F5, and F6 understood how the constraints were used to generate the recommended schedules, while F3–F6 felt that the recommended schedules aligned with their constraints. F3 and F4, however, did not initially notice that the tool was selecting a subset of their available hours in its recommendations, likely due to the more restricted availability that they had entered.
Reliance and Compliance
The extent to which participants relied on and complied with the tool was also variable. In particular, we did not observe any consistency in patterns of behaviour even within the same condition. For reliance, F2 and F3 relied on the tool’s recommendations to the extent that they actively scheduled their day around reviewing the daily schedules; F5 and F6 viewed the tool more as a reference that they sometimes incorporated into their planning; and F1, F4, and F7 did not rely on the tool’s recommendations much. For compliance, F1–F3, F5, and F6 reported having adjusted their schedules in response to the recommendations at some points (with F5 in particular striving to perfectly follow the recommendations on weekdays but not on weekends), while F1, F2, F4, and F6 also reported instances where they maintained their existing schedules over the tool’s recommendations.
F1–F6 all reported that the extent to which they were motivated to comply with the recommended schedule depended upon how the tool’s estimates compared to their expectations or actual earnings. For instance, F3 decided to stop following the schedule on what she felt was a “slow day”; F4 was intrigued by optimistic hourly estimates that were much higher than his personal goal for those times, despite mostly choosing not to follow the schedule; F5 became more motivated to follow the schedules after observing an increase in earnings of 20–30%; and F6 chose not to follow the schedule for hours where he felt that the tool was underestimating potential earnings. In terms of longitudinal effects, F1, F2, and F5 indicated that the outcomes of their first one or two days of interaction impacted their willingness to follow the recommendations for the rest of the study, again in line with the findings of (Mou and Cohen, 2017; Mou et al., 2017). Meanwhile, F4 stated that he lost trust in the system after seeing that its estimates did not align with his expectations.
Perceptions of Uncertainty
One of the causes for participants selectively complying with the recommended schedules at some times but not others was the variability in the accuracy of its estimates. F1–F6 mentioned cases where the hourly earning estimates fell below their expected or actual earnings; F1–F4 and F6 mentioned cases where they fell above; and F5 and F6 mentioned cases where they were similar. However, this variability was less evident in daily and weekly estimates: in Condition 2, F1, F2, and F6 felt that their daily earnings were accurately estimated, and F1 and F2 also felt that their weekly earnings were accurately estimated. Regardless, F1, F3, and F5 indicated that they used the tool’s estimates as goals for their earnings. Some correctly recognised that external factors could impact the tool’s accuracy: for F1 and F3, their own behaviour; and for F3 and F6, inherent environmental uncertainty. F2 and F6 believed that the tool estimated their base earnings before bonuses and tips, whereas it actually estimated their total earnings.
Among participants in the conditions with ranges (Conditions 3 and 4), F3–F5 all liked its presence, although F7 did not. F3–F5 also all reported focusing on the optimistic estimate over the mean estimate to some extent. However, a major cause of F4 losing trust was that he considered the pessimistic estimates to be too low; F5 echoed this sentiment but otherwise had a positive impression. When prompted regarding whether they thought a driver other than themselves (e.g., F7) could have earned similarly to the pessimistic estimate, both F4 and F5 found it difficult to fathom, despite the hedging text presented to F4 by Condition 4. Among participants in the conditions without ranges (Condition 2), F1 and F2 indicated that they would have appreciated the range, while F6 suggested that the range would have had limited value since he was already viewing the mean estimates as being potentially pessimistic.
Suggestions for Improvement
Participants suggested a number of features that could improve the tool’s utility and trustworthiness. These included the ability to understand how bonuses, tips (F2, F6), and events (F5) affect earnings; to estimate driver supply (F2) and trip mileage (F3); to restrict or view estimates for different regions of their cities (F1, F6); to incorporate pre-scheduled gigs into the recommended schedule (F6); and to be able to integrate the tool more smoothly with gig platforms, other assistants such as Gridwise, or personal calendar apps (F1, F5, F6).
7. Limitations
Our central motivation in this work was to assess the effects of different framings of uncertainty on trust and reliance in an AI decision aid. To this end, we attempted to isolate the effect of this design dimension from others by first iterating on our tool’s design. In particular, we used two separate pilots to identify and implement a number of design desiderata based on participant feedback. However, we certainly cannot claim that the design of our tool was optimal for engendering trust, as also reflected in feedback from our interview participants (Section 6.2.2). Data availability limitations, for instance, meant that our tool was unable to provide more direct feedback on how participants’ performance compared to the tool’s estimates on a daily basis. We cannot exclude the possibility that some of these design choices had compounding impacts on our measurements of trust and reliance.
Also, despite our best efforts, we were only able to recruit a small sample of drivers from the four cities and six platforms for which we could access historical data. These individuals were at least aware, if not active users, of the Gridwise app, and thus they may have been more focused on the outcomes of their driving compared to the general gig driver population. One benefit of recruiting from this population is that we received high-quality responses and experienced relatively low attrition, which facilitated our analysis. However, we envision that a future study aimed at a broader population could uncover an additional wealth of insights, especially considering the variation that existed even in our small sample.

### v2

@@ 6. Qualitative Analysis
P1: A 47-year-old male with a high school degree driving for Instacart and Lyft in Los Angeles, (D)
@@
P2: A 42-year-old female with a high school degree driving for Uber Eats in Los Angeles, (D)
@@
P3: A 29-year-old female with a graduate degree driving for Lyft in Chicago, (RH)
@@
P4: A 49-year-old male with an undergraduate degree driving for Lyft and Uber in Los Angeles, (RH)
@@
P5: A 46-year-old male with an undergraduate degree driving for DoorDash, GrubHub, and Uber Eats in Chicago, (R)
@@
P6: A 39-year-old male with a graduate degree driving for Lyft, Uber, and Uber Eats in Los Angeles, (D)
@@
P7: A 39-year-old male with a professional degree delivering for DoorDash in Los Angeles by bike, (RH)
Transcripts for the interviews were generated by Zoom. The author who conducted the interviews reviewed and corrected these transcripts, then performed structural coding (Saldaña, 2015a). Afterwards, two authors separately used QualCoder 3.3 (Curtain, 2023) to perform open coding (Saldaña, 2015a) and axial coding (Saldaña, 2015b). The two authors met to reconcile their codes and construct a unified codebook. Finally, the interviewing author re-applied the updated codes. Now, we discuss our findings in relation to participants’ motivations for using the tool (Section 6.1), perceptions of its accuracy (Section 6.2), and perceptions of its uncertainty based on the design conditions (Section 6.3).
6.1. Motivations and Routines
Participants reported a diversity of driving motivations and routines, which impacted their perceptions of the schedule recommendation tool’s usefulness. P1 and P2 viewed gig driving as a primary source of income, and thus found more value in the tool’s earnings estimates: “[The tool was] definitely worthwhile, just because it gave me a number, a projection. […] They definitely motivate me to keep going the next day.” – P1 (D) Meanwhile, P3–P7 used gig driving to supplement other sources of income, but P4 still viewed his earning goals as important. Unlike other participants, P7 delivered with a bicycle in his spare time. He felt that his current commitment was insufficient to want to use the tool more: “If I take this job to a full time, take it seriously? I would [want to use it more].” – P7 (RH) Nevertheless, drivers found value in the tool regardless of their level of motivation. P1–P6 all reported challenges in estimating their potential earnings as a consequence of unpredictability in gig demand, pay, or location, or of gig platforms providing insufficient information. For instance: “Uber’s details that they offer to drivers through their interfaces are sorely lacking. So I’m grateful for the opportunity to interact with this tool.” – P6 (D)
6.2. RQ1: Perceptions of Accuracy
When evaluating the tool’s accuracy, participants weighed its recommendations against their own routines and intuitions. For P1, P3, P6, and P7, the tool was a reference for how they could perform in their existing routines, rather than something that would reshape their routines: “I still would’ve followed my routine. […] I was fortunate enough to at least have the tool make me a schedule based on the routine that I currently do.” – P3 (RH) However, P5 suggested that the tool could use a question-answering approach to nudge users into altering their routines, by first understanding their activity patterns and then suggesting modifications. When the tool was inaccurate, participants reacted in different ways. P1, P2, and P4 observed that instances of the tool being inaccurate decreased their desire to comply with the tool’s recommendations: “If I was making more than what it said, I would have done it more consistently on the schedule.” – P2 (D) P4’s reactions to inaccuracies were influenced by his expectations. He was motivated on one instance by the tool’s estimates exceeding his goals: “My target’s […] $30 an hour. Because those levels were consistently below $30, […] I wasn’t motivated to study it. But when I saw the 4 to 6 am, that kind of piqued my interest.” – P4 (RH) Maintaining consistent perceptions of accuracy over time was important for building trust in this context. P1, P2, and P5 indicated that the outcomes of their first one or two days of interaction impacted their willingness to follow the recommendations for the rest of the study. P4 and P5 indicated that their use of the tool would be strengthened longitudinally if they consistently perceived its predictions as being accurate: “Once I learned that it was accurate, and I had trust in it, and it was really helping, then I’d probably use it more and more.” – P4 (RH)
6.3. RQ2: Perceptions of Uncertainty
Some participants recognised that the accuracy of the tool’s earnings estimates would be impacted by both their own decisions (P1, P3) and other environmental factors (P6): “It also depends, too, on the rides that I accept.” – P3 (RH) “It might be true that I might earn the forecasted average earnings. But surges can definitely make a difference.” – P6 (D) Note that the tool’s uncertainty was not exposed to P6, suggesting that this observation originated from their innate mental model. Recognising the effect of their own agency led P1 and P3, as well as P5, to adopt the tool’s estimates as goals for their own earnings: “[…] setting daily goals of how much money I would like to make […] was definitely something that I wasn’t really doing prior to doing this study or using this tool.” – P3 (RH) Also, P1 suggested that being able to compare their earnings to the tool’s estimates in an hourly breakdown would be helpful for goal-setting.
All participants in Conditions (R) and (RH) (P3–P5, P7) appreciated the presence of ranges. P5 compared the tool’s range to his own experiences: “I was always over the average. So, to me, I was kind of in my head using that as a low.” – P5 (R) For participants in the other conditions, P1 and P2 indicated that they would’ve preferred to have had ranges. However, P6 suggested that ranges may lead to disappointment when they are used for goal-setting: “If they earn less than [the higher number], then they probably might feel disappointed in the tool through no fault of its own, right? If you say $12 to $18, and it comes in at $14, […] I could understand how folks might look at that as a let down.” – P6 (D) Thus, ranged-based uncertainty was useful for decision-making, but needed to be calibrated against expectations.
Both P4 and P5 struggled with the idea that the uncertainty in the tool’s estimates could have originated from drivers with habits different to themselves: “Obviously no one’s ever gonna work if it’s just $4 or $5 an hour.” – P4 (RH) This was in spite of the lexical hedging presented to P4. At least for this participant, the verbiage in our hedges may thus have failed to achieve its goal of leading him to consider potential sources of uncertainty more carefully.
7. Discussion
7.1. Key Findings and Implications
Trust in AI decision aids is built both initially and over time. We found that participants’ pre-interaction trust in our schedule recommendation tool significantly impacted their trust during interaction (Sections 5.2 and 5.3.1). This is consistent with findings in the medical domain that practitioners (Burgess et al., 2023; Cai et al., 2021) and patients (Mou and Cohen, 2017; Mou et al., 2017) prefer to gauge their trust prior to interaction. Our interviews similarly showed that perceptions of the tool’s accuracy in the first two days influenced subsequent trust (Section 6.2). Yet, we also found that trust and reliance increased across interactions with the tool. While perceived accuracy had diminishing impacts on trust in later stages of the user study (Section 5.2), P4’s experience (Section 6.2) shows that critical incidents where estimates differ significantly from expectations can cause catastrophic losses of trust (Stanley and Dorton, 2023). P4 found this difficult to recover from. However, losses of trust could be mitigated prospectively by calibrating perceptions of accuracy, e.g. by emphasising that drivers’ outcomes are also a function of their own decisions. This could foster appropriate reliance by helping users decide when or when not to rely on the tool (Schemmer et al., 2023).
Interactivity could help to maintain trust over time. Losses of trust can also be mitigated retrospectively based on trust repair strategies (Pareek et al., 2024). Based on our qualitative findings, we hypothesise two mechanisms by which interactivity could help to maintain and repair trust. First, interactivity may enhance perceptions of control. The modes of interaction suggested by our participants, such as hourly breakdowns and question-answering, would assure users that the AI has the intent and agency to capture and learn from their preferences (Pareek et al., 2024). Second, interactivity could help users to better recall their experiences and decisions. Most interview participants could not recall whether their earnings significantly differed from the tool’s estimates until the interviewer probed further.
The impact of exposing uncertainty on trust in AI decision aids depends on task alignment. Prior work has reached mixed conclusions on how exposing uncertainty in AI impacts trust and reliance. On similar tasks, Zhang et al. (2020) found that confidence scores improve reliance, whereas Prabhudesai et al. (2023) found that distribution plots dampen trust and reliance. Yang et al. (2024) found that the effects of these designs depended on individual characteristics, but our results suggest another dimension: task alignment in the designs themselves. Task-aligned uncertainty representations, i.e. scalar ranges as opposed to distributions, allowed our participants to incorporate uncertainty directly into their decision-making (Section 6.3), thus improving trust (Section 5.3.2). This is consistent with findings in the AI explainability literature that domain-aligned explanations are more persuasive (Cau et al., 2023; Naiseh et al., 2023). We hypothesise that task alignment also underlies the negative effect of hedging we observed (Section 5.3.2): thinking about other drivers is not helpful when drivers are trying to reason about their own outcomes (Section 6.3).
How uncertainty is exposed should be adapted to user subpopulations. Our results did not find that a one-size-fits-all approach exists to fostering trust. Even within the same condition, participants exhibited variability in how they reacted to the outcomes of their reliance. Nevertheless, we hypothesise that it may be helpful to adapt uncertainty in predictions and recommendations to subgroups within the gig driver population. Specifically, our qualitative results point to differing perceptions of accuracy and uncertainty between highly motivated drivers (e.g. P1 and P2) and less motivated drivers (e.g. P7). If drivers received estimates from those with habits similar to themselves, this could mitigate P4 and P5’s struggles with how to interpret uncertainty. A future large-scale study could help to confirm our hypothesis.
7.2. Limitations
Our work has two primary limitations. First, we cannot claim that the design of our tool was optimal for engendering trust. Our focus was on testing how designs for exposing uncertainty would impact trust and reliance. Thus, we attempted to isolate the effect of this design dimension by refining the tool through a formative study and pilot (Appendix B). Nevertheless, further improvements may have been possible. For instance, we could not provide retrospective breakdowns of participants’ earnings due to data availability limitations. Thus, design choices orthogonal to the exposure of uncertainty may have impacted participants’ trust and reliance.
Second, despite our best efforts, our sample of drivers was limited. These individuals were at least aware, if not active users, of the Gridwise app, and thus they may have been more focused on their outcomes than the general gig driver population. The trust of users in an AI decision aid is contingent upon their domain knowledge (Kim et al., 2023; Wysocki et al., 2023), and — as we demonstrate (Section 6.1) — the extent to which they integrate the decision aid into their existing routines. A future study aimed at a broader population of gig drivers could uncover additional insights by explicitly controlling for factors such as full-time status and driver tenure. We were also unable to reach participants who discontinued their participation. Future studies that follow up with such participants would be a valuable source of data on mechanisms of trust loss and repair in longitudinal settings.
7.3. Recommendations for Future Work
The paucity of similar longitudinal, in-situ studies in prior work is understandable given the logistical challenges we encountered. We stress the importance of observational studies to improve domain understanding as a basis for longitudinal interventional studies. Our pilot interviews helped us to design a tool that was task-aligned, which led participants to find value in it over repeated interactions. Furthermore, our study design aimed to increase perceived control while reducing user burden through flexibility in the scheduling of participation; customisability of the constraints on the tool; shorter survey instruments; and incremental compensation.